OpenStudy (anonymous):

What is F(3)? If your answer is not a whole number, enter it as a fraction in lowest terms, using the slash mare (/) as the fraction bar. F(x)=x^2+5/4x-3

jigglypuff314 (jigglypuff314):

you would plug in 3 for x so f(3) = (3)^2 + 5/4 (3) - 3

OpenStudy (anonymous):

So I would get f(3)=14/9? What after that?

jigglypuff314 (jigglypuff314):

not quite :) f(3) = (3)^2 + 5/4 (3) - 3 = (3)(3) + 5/4 (3) - 3 = 9 + 15/4 - 3 = 6 + 15/4 = 24/4 + 15/4 = 39/4 therefore f(3) = 39/4

OpenStudy (anonymous):

Thank you so much! So in lowest terms that would be 9 and 3/4?

jigglypuff314 (jigglypuff314):

if you want it as a mixed fraction then yes :) otherwise 39/4 is just fine.

OpenStudy (anonymous):

Thank you :D so much!

jigglypuff314 (jigglypuff314):

you're welcome :)